OpenStudy (anonymous):

Find the linearization L(x) of the function f(x)=x^3/4 at x=256.

OpenStudy (anonymous):

i got (3/16)x-16. the computer said it was wrong

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

sorry 64+(3/16)(x-256)

OpenStudy (phi):

maybe they want you to simplify that

OpenStudy (anonymous):

i did i got (3/16)x-16

OpenStudy (phi):

how did you get -16 ?

OpenStudy (anonymous):

when you expand 64+(3/16)(x-256 you get 64+(3/16)x-48 right?

OpenStudy (phi):

yes, and 64-48 = +16

OpenStudy (anonymous):

never mind its 16

OpenStudy (anonymous):

oh my i am really failing today

OpenStudy (phi):

The good news is you know how to do the hard part.

OpenStudy (anonymous):

haha true. thanks

OpenStudy (phi):

fixing the easy stuff is not that hard. Just slow down, and double check your arithmetic

OpenStudy (anonymous):

i have another question. its telling me to use linear approximation to approximate (8.07)^(2/3). i got 4.023. it says i am wrong though

OpenStudy (phi):

f(x) = x^(⅔) f'(x) = (⅔) x ^(-⅓) expand around x0= 8 f(8)= 4 f'(8) = ⅔ * ½= ⅓ f(x)= f(x0) + f'(x0)* (x-x0) f(x) = 4 + ⅓ * (x-8) f(8.07)= 4+ ⅓(.07)= 8.02333... my calculator says 8^(⅔) is 4.023299437

OpenStudy (anonymous):

yeah none of that works

OpenStudy (phi):

did you try more digits ? 4.02333

OpenStudy (anonymous):

yup

OpenStudy (phi):

See if they have any instructions on how many digits they want. If you can't find that, make a note and ask your teacher, because the answer is 4.023 (3 repeating)

OpenStudy (anonymous):

this is the question "Use a linear approximation to estimate the number (8.07)^2/3."

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

there is an outside chance there is a typo, and they meant 8.075 (because 0.075 is evenly divisible by 3) you would get 4.025

OpenStudy (anonymous):

nope its not that. do you think you could help me out on one last problem

OpenStudy (phi):

and just to be sure, ⅔ is the exponent, and it is not (8.07^2 ) * 1/3

OpenStudy (anonymous):

its 2/3

OpenStudy (anonymous):

Find the linearization L(x) of the function g(x)=xf(x2) at x=2 given the following information. f(2)=−1 f′(2)=8 f(4)=4 f′(4)=−2

OpenStudy (phi):

is that \[ g(x) = x f(x^2) \] ? if so, use an ^ to show an exponent

OpenStudy (anonymous):

yeah its that. my bad

OpenStudy (phi):

find the usual stuff: g(2) and g'(2) what do you get ?

OpenStudy (anonymous):

g(2)=8 but i dont know how to find g'(x). would g'(x)=f(x^2)+2x^2f(x^2)?

OpenStudy (phi):

product rule: g(x) = x f(x^2) g' = x f'(x^2) * 2x + f(x^2) or g'= f(x^2) + 2 x^2 f'(x^2) yes, it looks the same as what you got. now evaluate at x=2

OpenStudy (anonymous):

it would be 36

OpenStudy (phi):

how?

OpenStudy (anonymous):

4+8(4)=4+32=36

OpenStudy (phi):

g'= f(x^2) + 2 x^2 f'(x^2) f(4) + 2* 2^2 * f'(4)

OpenStudy (anonymous):

=36

OpenStudy (phi):

f(2)=−1 f′(2)=8 f(4)=4 f′(4)=−2 isn't f'(4) = -2 ?

OpenStudy (anonymous):

oh my. i cant get anything right. so it would be -12

OpenStudy (anonymous):

FINALLY!!!!! i got it its -12x+32. thanks so much

OpenStudy (phi):

Excellent!